OpenStudy (anonymous):

Jaylen has decided to purchase a $14,750 car and would like to finance it for five years. Her bank has offered her a standard 6.9% APR, while the dealership has offered her a special 3.9% APR. How much money will she save each month by choosing financing through the dealership? A. $29.45 B. $24.61 C. $20.39 D. $43.25

OpenStudy (wolf1728):

You will need the Loan Payment Formula: http://www.1728.org/loanform.htm MnthPmt = rate + [rate/((1+rate)^months)-1) ] * principal WHERE RATE IS THE MONTHLY RATE SO RATE = .069/12 = MnthPmt = 0.00575 + [.00575/((1.00575^60)-1)*14,750 MnthPmt = (0.00575 + (0.00575/1.41059543682032000)) * 14,750 MnthPmt = (0.00575 + (0.00575/.41059543682032000)) * 14,750 MnthPmt = (0.00575 + 0.0140040523697204)*14,750 MnthPmt = (0.0197540523697204)*14,750 MnthPmt = 291.37 FOR RATE = 3.9 MONTHLY RATE = .00325 MnthPmt = 0.00325 + [.00325/((1.00325^60)-1)*14,750 MnthPmt = 0.00325 + [.00325/((1.00325^60)-1)*14,750 MnthPmt = 0.00325 + [.00325/(1.21492677792396-1)]*14,750 MnthPmt = 0.00325 + [.00325/(.21492677792396)]*14,750 MnthPmt = [0.00325 + 0.0151214289414874]*14,750 MnthPmt = 0.0183714289414874 * 14,750 MnthPmt = 270.98 291.37 - 270.98 = 20.39 Difference Answer is C